OpenStudy (anonymous):

George jogged a total distance of 9 and 1 over 2 miles during the months of October and November. If George only jogged 1 over 4 mile every day, which expression shows the number of days in which he went jogging? 9 and 1 over 2 multiplied by 1 over 4 9 and 1 over 2 plus 1 over 4 9 and 1 over 2 divided by 1 over 4 9 and 1 over 2 minus 1 over 4

OpenStudy (anonymous):

help ill give fan and medal

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (mathstudent55):

Let's say he jogged 1 mile per day. Then in 5 days, he would have jogged 5 miles. If you divide 5 miles (the total distance jogged) by 1 mile (the distance jogged in one day), you would get 5/1 = 5 which is the number of days we said he jogged. This shows you that you need to divide the total number of miles jogged by the number of miles jogged in each day to find the number of days.

OpenStudy (anonymous):

so 9 divided by 2

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Notice what the numbers are: 9 and 1 over 2 is really the mixed numeral \(9 \dfrac{1}{2} \), or 9.5 or nine and a half. 1 over 4 is the fraction \(\dfrac{1}{4} \), or 0.25 or one quarter. You need to divide \(9 \dfrac{1}{2} \) by \(\dfrac{1}{4} \), so the answer is choice C.

OpenStudy (triciaal):

total jogged = amount per day * number of days 9 and 1/2 = 1/4 * d, where d is the number of days solve for d